# How informative are summaries of the cosmic 21-cm signal?

### v1

@@ 5.2 Comparison of all summaries
Finally, our concatenation of the 2DPS and IMNN significantly outperforms all of the individual summaries (c.f. the red curve in Fig. 6). As we saw above, these two summary statistics are highly complimentary: the 2DPS is robust (i.e. a narrow information PDF), while the IMNN summary can be much more constraining for some parameters but much less for others (i.e. it has a very broad information PDF). The distribution of Fisher information of 2DPS + IMNN retains the narrowness of the 2DPS, but is shifted to the right, including the high-value tail from the IMNN.  [^13] We note that in the case in which some value of the IMNN summary would be completely degenerate with some 2DPS bin, the resulting covariance matrix would be singular and therefore non-invertible. This would make estimating the Fisher matrix unreliable. However, as the IMNN compression and 2DPS are very different summaries, this is unlikely to be a problem. The combined summary is a factor of $\sim$6.5–9.5 better in the single parameter variance (Eq. 26), averaged over the prior samples. Combining multiple 21-cm statistics has already been proven to tighten parameter constraints (e.g. Gazagnes et al. 2021; Watkinson et al. 2022). However, previous work using traditional inference was limited by the need to derive a tractable likelihood which becomes impossible for non-trivial summaries. Fortunately, recent advances in applying SBI to 21-cm (e.g. Zhao et al. 2022a, b; Prelogović & Mesinger 2023; Saxena et al. 2023) will allow us to cleanly combine multiple summary statistics without having to explicitly define the likelihood. Unfortunately, the neural density estimation that is intrinsic to SBI can be unstable and/or require large training sets, if the summary statistics become very high dimensional. This work illustrates how combining only two complimentary statistics can result in much tighter, more robust parameter constraints than would be available using either individually.
5.3 One-term vs. full Fisher matrix
@@
Figure 7: Distributions of Fisher information for the 1DPS (red) and 2DPS (blue). Solid lines account for both terms in the Fisher matrix (c.f. Eq. 4), while the dashed lines correspond to the common simplification of a constant covariance (i.e. including only the first term in Eq. 4). The assumption of a constant covariance underpredicts the Fisher information by a factor of $\sim 2$. For a 5-parameter model, this translates to a $\sim 15\%$ larger single parameter variance (Eq. 26).
Figure 8: Analogous to Fig. 7, but for the other 21-cm summary statistics.
@@
The full Fisher matrix of a Gaussian likelihood can be separated in two terms $\bm{F}=\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}}$ (see Eq. 4), where the first term includes the gradient of the mean and the second term the gradient of the covariance matrix (e.g. Appendix B; Tegmark et al. 1997). All of our previous results were calculated using the full expression, including both terms. However, it is common in the literature to ignore one of the terms, by approximating the mean or the covariance matrix as a constant (e.g. Carron 2013). In most of cosmological Fisher forecasts using the power spectrum (e.g. Sailer et al. 2021; Abazajian et al. 2022; Euclid Collaboration et al. 2022; Mason et al. 2023; Bykov et al. 2023), the second term, $\bm{F}_{\bm{\Sigma}}$, is neglected. [^14] For examples of ignoring the first term, see e.g. Abramo (2012); d’Assignies D et al. (2023). In some cases, the covariance matrix can indeed be independent or only weakly dependent on the parameters, justifying ignoring the second Fisher term. However, the importance of the $\bm{F}_{\bm{\Sigma}}$ term for 21-cm Fisher forecasts has not yet been investigated (e.g. Greig et al. 2022; Balu et al. 2023; Mason et al. 2023; Hothi et al. 2023). Indeed in Prelogović & Mesinger (2023) we showed that assuming a parameter-independent covariance matrix of the 21-cm 1DPS likelihood can bias the posterior estimation. In this section, we quantify the importance of the $\bm{F}_{\bm{\Sigma}}$ term in determining the Fisher information [^15] Assuming a constant covariance matrix can result in other biases at the level of the full Fisher matrix. For example, it could impact parameter degeneracies and/or improve constraints of some parameters at the cost of others. Here we only focus on the effect this simplification has on the determinant of the Fisher matrix as this quantity measures the average parameter constraining power of the summary statistics., for all of our 21-cm summaries. It is important to note that adding this term will always increase the information content, as $\det(\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}})\geq\det\bm{F}_{\bm{\mu}}$. For more details, see the end of Appendix A.
@@
In Fig. 7 we plot the Fisher information PDF of the commonly-used 1DPS and 2DPS. The solid curves are the same as in the previous figure, corresponding to the full expression for the Fisher information. The dashed curves only include the first term, assuming a constant covariance matrix. For both 1D and 2D power spectra, neglecting the second term underestimates the Fisher information by a factor of $\sim 2.3$ on average over the prior samples. For our 5-parameter model, this translates into an underestimate of the single parameter variance (Eq. 26) by 15%.
@@
In Fig. 8 we compare the Fisher information PDFs with and without the second Fisher term for the remaining individual summaries. We see that for the other 21-cm summaries, the second Fisher term is even more important than for the power spectra. For example, assuming a constant covariance underestimates the mean Fisher information of the RNN for a factor of $\sim$4700 and a factor of $\sim$11100 for the wavelets, averaged over the prior samples. Note that although the IMNN is trained to normalize the covariance to unity, thus ensuring a constant covariance matrix by construction, this is only done at $\bm{\theta}_{\text{fid}}$. For the other points in the parameter space, there is no guarantee that this condition is met.
@@ 6 Conclusions
The cosmic 21-cm signal will revolutionazie our understanding of the CD and Eo). ”Big Data” analysis techniques will be required to cope with radiotelescopes such as the Square Kilometre Array (SKA). In this work we compare several approaches of compressing 21-cm lightcone data, on the basis of their Fisher information. Specifically, we compare the determinant of the Fisher information for the following summary statistics: (i) the spherically-averaged power spectrum (1DPS), (ii) the cylindrically-averaged power spectrum (2DPS), (iii) the 2D Wavelet scattering transform (WST), (iv) a recurrent neural network (RNN) trained as a regressor; (v) an information-maximizing neural network (IMNN); and (vi) the combination of 2DPS and IMNN. Importantly we compare their Fisher information across parameter space, since there is currently no standard model for the astrophysics of the first galaxies, which determine the cosmic 21-cm signal.
@@
The 2DPS is the individual summary with the highest median Fisher information (Fig. 6, Table 1). The distribution of its Fisher information is also relatively narrow across parameter space. This means that compared to most other summaries, the 2DPS can recover astrophysical parameters with relatively similar precision regardless of what are the ”true” values of our Universe. We also show that the 2DPS slightly outperforms the 1DPS, confirming that the anisotropy of the 21-cm signal helps constrain parameters. The information content of both RNN and wavelets is generally lower than either the 1DPS or the 2DPS. However, we caution that both of these summaries could be defined in several different ways, potentially improving their performances (e.g. Greig et al. 2022, 2023; Hothi et al. 2023; Zhao et al. 2023a).
@@
In this work, we also introduce IMNNs to the field of 21-cm cosmology. Although capable of achieving the highest Fisher information for some parameter choices, the IMNN does not generalize well, resulting in a broad distribution across the prior volume. Indeed, we find an enormous difference between its Fisher information computed at the parameter set used for training, $\bm{\theta}_{\text{fid}}$, and its information computed at the remaining prior samples (Fig. 5). This means that the IMNN overfits to the information content of the fiducial. The regularization inherent to the IMNN loss (c.f. Eq. 22 and the discussion in Footnote 8) makes it difficult to avoid this overfitting by training over many different choices of $\bm{\theta}_{\rm fid}$. A simpler solution could be to first perform inference with a less optimal summary, in order to determine a maximum likelihood point, and then train the IMNN at that point.
@@
Combining the IMNN with the 2DPS results in a summary whose median Fisher information is almost $\sim$4 orders of magnitude larger than the 2DPS alone. This results in a factor of $\sim$6.5–9.5 tighter single parameter variance averaged over the prior samples. This means the IMNN extracts complementary information with respect to the 2DPS. Performing simulation-based inference using such a summary could yield extremely tight and robust parameter constraints. Moreover, combining only these two complementary summaries still preserves a high level of compression, with our lightcone of $\sim 1.6\cdot 10^{6}$ cells reduced to just 77 numbers.
@@
Finally, we stress the importance of not assuming a constant covariance when performing Fisher forecasts using 21-cm observables. This common assumption underestimates the Fisher information of the power spectrum on average by a factor of $\sim$2.3, corresponding to a 15% increase in the single parameter variance (Eq. 26). For other summaries, the assumption underestimates the Fisher information by orders of magnitude.
@@ Acknowledgements.
D.P. thanks T. Charnock for many discussions and in depth explanations on how the IMNN function, as well as L. Makinen on help with stabilizing IMNN training. A.M. acknowledges support from the Ministry of Universities and Research (MUR) through the PRIN project ”Optimal inference from radio images of the epoch of reionization” as well as the PNRR project ”Centro Nazionale di Ricerca in High Performance Computing, Big Data e Quantum Computing”. We gratefully acknowledge computational resources of the HPC center at SNS. We acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support (grant IMC21cm - HP10CWIEF7). This work used the Extreme Science and Engineering Discovery Environment (XSEDE), which is supported by National Science Foundation grant number ACI-1548562. Specifically, it used the Bridges-2 system, which is supported by NSF award number ACI-1928147, at the Pittsburgh Supercomputing Center (PSC).
@@ Appendix A Proof of the determinant inequality
Let us consider two matrices $\bm{C}$ and $\bm{F}^{-1}$ of shape $m\times m$ for which the following inequality holds
@@
From this, it follows:
@@
$\det\bm{C}$ $\geq\det\bm{F}^{-1}$ (33)
Here the trick was to push the inequality towards the multivariate Gaussian integral. After integrating, what is left is only the inverse of a normalization constant. Notice that if one wants to interpret $\bm{C}$ as a covariance matrix and $\bm{F}$ as a Fisher matrix, then it would make more sense to integrate over $\bm{x}^{T}\bm{C}^{-1}\bm{x}$ and $\bm{x}^{T}\bm{F}\bm{x}$. However this is irrelevant for the proof, as it would only result in the change of the inequality direction in each row, ending with $\det\bm{C}^{-1}\leq\det\bm{F}$. We usually assume the matrices are non-singular and thus invertible, which is needed here.
Note that the identical procedure can be used to prove that the one-term Fisher is a lower bound of the two-term Fisher. For $\bm{F}=\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}}$ one can write $\bm{x}^{T}(\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}})\bm{x}\geq\bm{x}^{T}\bm{F}_{\bm{\mu}}\bm{x}$, which by following the same steps translates into $\det(\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}})\geq\det\bm{F}_{\bm{\mu}}$.
@@ Appendix B Derivation of a Gaussian Fisher information matrix
In order to derive Eq. 4, i.e. the Fisher information of a multivariate Gaussian distribution, we start by writing the Fisher information matrix equation (Eq. 1) in the equivalent form:
@@
The first one follows from $\bm{\Sigma}\bm{\Sigma}^{-1}=\mathbb{1}$ and calculating its derivative, the second one is Jacobi’s relation for the derivative of the determinant. The last two are expectation relations for Gaussian distributions for an arbitrary matrix $\bm{A}$ (where $\dim\bm{A}=\dim\bm{\Sigma}$). The first of these follows from the anti-symmetry of the integral, and the second is left as an exercise for the conscientious reader.
@@
and computing its second derivative, we get:
@@
$\phantom{=}-\left(\partial_{m}\bm{\mu}^{T}\bm{\Sigma}^{-1}\partial_{n}\bm{\Sigma}+\partial_{n}\bm{\mu}^{T}\bm{\Sigma}^{-1}\partial_{m}\bm{\Sigma}\right)\bm{\Sigma}^{-1}(\bm{d}-\bm{\mu})$ (48)
Taking the negative expectation value of the final equation can now be done by using relations 39 and 40. One can easily see that only the first three terms contribute, with others canceling each other or being equal to zero. With this, we finally have:

### v2

@@ 5.2 Comparison of all summaries
Finally, our concatenation of the 2DPS and IMNN significantly outperforms all of the individual summaries (c.f. the red curve in Fig. 7). As we saw above, these two summary statistics are highly complimentary: the 2DPS is robust (i.e., a narrow information PDF), while the IMNN summary can be much more constraining for some parameters but much less for others (i.e., it has a very broad information PDF). The distribution of Fisher information of 2DPS + IMNN retains the narrowness of the 2DPS, but is shifted to the right, including the high-value tail from the IMNN.  [^16] We note that in the case in which some value of the IMNN summary would be completely degenerate with some 2DPS bin, the resulting covariance matrix would be singular and therefore non-invertible. This would make estimating the Fisher matrix unreliable. However, as the IMNN compression and 2DPS are very different summaries, this is unlikely to be a problem. The combined summary is a factor of $\sim$6.5–9.5 better in the single parameter variance (Eq. 26), averaged over the prior samples. Combining multiple 21 cm statistics has already been proven to tighten parameter constraints (e.g., Gazagnes et al. 2021; Watkinson et al. 2022). However, previous work using traditional inference was limited by the need to derive a tractable likelihood, which becomes impossible for nontrivial summaries. Fortunately, recent advances in applying SBI to 21 cm (e.g.,Zhao et al. 2022a, b; Prelogović & Mesinger 2023; Saxena et al. 2023) will allow us to cleanly combine multiple summary statistics without having to explicitly define the likelihood. Unfortunately, the neural density estimation that is intrinsic to SBI can be unstable and/or require large training sets, if the summary statistics become very high dimensional. This work illustrates how combining only two complimentary statistics can result in much tighter, more robust parameter constraints than would be available using either individually.
5.3 One-term versus full Fisher matrix
@@
Figure 8: Distributions of Fisher information for the 1DPS (red) and 2DPS (blue). Solid lines account for both terms in the Fisher matrix (c.f. Eq. 4), while the dashed lines correspond to the common simplification of a constant covariance (i.e., including only the first term in Eq. 4). The assumption of a constant covariance underpredicts the Fisher information by a factor of $\sim 2$. For a 5-parameter model, this translates to a $\sim 15\%$ larger single parameter variance (Eq. 26).
Figure 9: Analogous to Fig. 8, but for the other 21 cm summary statistics.
@@
The full Fisher matrix of a Gaussian likelihood can be separated in two terms $\bm{F}=\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}}$ (see Eq. 4), where the first term includes the gradient of the mean and the second term the gradient of the covariance matrix (e.g., Appendix B; Tegmark et al. 1997). All of our results from the previous sections were calculated using the full expression, including both terms. However, it is common in the literature to ignore one of the terms, by approximating the mean or the covariance matrix as a constant (e.g., Carron 2013). In most of cosmological Fisher forecasts using the PS (e.g., Sailer et al. 2021; Abazajian et al. 2022; Euclid Collaboration et al. 2022; Mason et al. 2023; Bykov et al. 2023), the second term, $\bm{F}_{\bm{\Sigma}}$, is neglected. [^17] For examples of ignoring the first term, see, e.g., Abramo (2012) and d’Assignies D et al. (2023). In some cases, the covariance matrix can indeed be independent or only weakly dependent on the parameters, justifying ignoring the second Fisher term. However, the importance of the $\bm{F}_{\bm{\Sigma}}$ term for 21 cm Fisher forecasts has not yet been investigated (e.g., Greig et al. 2022; Balu et al. 2023; Mason et al. 2023; Hothi et al. 2023). Indeed in Prelogović & Mesinger (2023) we showed that assuming a parameter-independent covariance matrix of the 21 cm 1DPS likelihood can bias the posterior estimation. In this section, we quantify the importance of the $\bm{F}_{\bm{\Sigma}}$ term in determining the Fisher information [^18] Assuming a constant covariance matrix can result in other biases at the level of the full Fisher matrix. For example, it could impact parameter degeneracies and/or improve constraints of some parameters at the cost of others. Here we only focus on the effect this simplification has on the determinant of the Fisher matrix as this quantity measures the average parameter constraining power of the summary statistics., for all of our 21 cm summaries. It is important to note that adding this term will always increase the information content, as $\det(\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}})\geq\det\bm{F}_{\bm{\mu}}$. More details are given at the end of Appendix A.
@@
In Fig. 8 we plot the Fisher information PDF of the commonly used 1DPS and 2DPS. The solid curves are the same as in the previous figure, corresponding to the full expression for the Fisher information. The dashed curves only include the first term, assuming a constant covariance matrix. For both 1D and 2D power spectra, neglecting the second term underestimates the Fisher information by a factor of $\sim 2.3$ on average over the prior samples. For our 5-parameter model, this translates into an underestimate of the single parameter variance (Eq. 26) by 15%.
@@
In Fig. 9 we compare the Fisher information PDFs with and without the second Fisher term for the remaining individual summaries. We see that for the other 21 cm summaries, the second Fisher term is even more important than for the power spectra. For example, assuming a constant covariance underestimates the mean Fisher information of the RNN for a factor of $\sim$2000 and a factor of $\sim$11100 for the wavelets, averaged over the prior samples. We note that although the IMNN is trained to normalize the covariance to unity, thus ensuring a constant covariance matrix by construction, this is only done at $\bm{\theta}_{\text{fid}}$. For the other points in the parameter space, there is no guarantee that this condition is met.
@@ 6 Conclusions
The cosmic 21 cm signal will revolutionize our understanding of the CD and EoR. ”Big Data” analysis techniques will be required to cope with radio telescopes such as the SKA. In this work we compare several approaches to compressing 21 cm light cone data, on the basis of their Fisher information. Specifically, we compare the determinant of the Fisher information for the following summary statistics: (i) the 1DPS, (ii) the 2DPS, (iii) the 2D WST, (iv) a RNN trained as a regressor; (v) an IMNN; and (vi) the combination of 2DPS and IMNN. Importantly we compare their Fisher information across the parameter space since there is currently no standard model for the astrophysics of the first galaxies, which determines the cosmic 21 cm signal.
@@
The 2DPS is the individual summary with the highest median Fisher information (Fig. 7, Table 1). The distribution of its Fisher information is also relatively narrow across the parameter space. This means that compared to most other summaries, the 2DPS can recover astrophysical parameters with relatively similar precision regardless of what the ”true” values of our Universe are. We also show that the 2DPS slightly outperforms the 1DPS, confirming that the anisotropy of the 21 cm signal helps constrain parameters. The information content of both the RNN and wavelets is generally lower than either the 1DPS or the 2DPS. However, we caution that both of these summaries could be defined in several different ways, potentially improving their performances (e.g., Greig et al. 2022, 2023; Hothi et al. 2023; Zhao et al. 2023a). Moreover, our Fisher estimates for the RNN and wavelet summaries require a larger dataset to fully converge (see Appendix D).
@@
In this work we also introduce IMNNs to the field of 21 cm cosmology. Although capable of achieving the highest Fisher information for some parameter choices, the IMNN does not generalize well, resulting in a broad distribution across the prior volume. We find an enormous difference between its Fisher information computed at the parameter set used for training, $\bm{\theta}_{\text{fid}}$, and its information computed at the remaining prior samples (Fig. 6). This means that the IMNN overfits to the information content of the fiducial. The regularization inherent to the IMNN loss (see Eq. 22 and the following discussion) makes it difficult to avoid this overfitting by training over many different choices of $\bm{\theta}_{\rm fid}$. A simpler solution could be to first perform inference with a less optimal summary, in order to determine a maximum likelihood point, and then train the IMNN at that point.
@@
Combining the IMNN with the 2DPS results in a summary whose median Fisher information is almost $\sim$4 orders of magnitude larger than the 2DPS alone. This results in a factor of $\sim$6.5–9.5 tighter single parameter variance averaged over the prior samples. This means the IMNN extracts complementary information with respect to the 2DPS. Performing SBI using such a summary could yield extremely tight and robust parameter constraints. Moreover, combining only these two complementary summaries still preserves a high level of compression, with our light cone of $\sim 1.6\cdot 10^{6}$ cells reduced to just 77 numbers.
@@
Finally, we stress the importance of not assuming a constant covariance when performing Fisher forecasts using 21 cm observables. This common assumption underestimates the Fisher information of the PS on average by a factor of $\sim$2.3, corresponding to a 15% increase in the single parameter variance (Eq. 26). For other summaries, the assumption underestimates the Fisher information by orders of magnitude.
@@ Acknowledgements.
D.P. thanks T. Charnock for many discussions and in depth explanations on how the IMNN function, L. Makinen on help with stabilizing IMNN training, and C. Sui for the discussion on Mutual Information. We thank B. Greig and S. Murray for many useful discussions and for giving the comments on the manuscript. A.M. acknowledges support from the Ministry of Universities and Research (MUR) through the PRIN project ”Optimal inference from radio images of the epoch of reionization” as well as the PNRR project ”Centro Nazionale di Ricerca in High Performance Computing, Big Data e Quantum Computing”. We gratefully acknowledge computational resources of the HPC center at SNS. We acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support (grant IMC21cm - HP10CWIEF7). This work used the Extreme Science and Engineering Discovery Environment (XSEDE), which is supported by National Science Foundation grant number ACI-1548562. Specifically, it used the Bridges-2 system, which is supported by NSF award number ACI-1928147, at the Pittsburgh Supercomputing Center (PSC).
@@ Appendix A Proof of the determinant inequality
We considered two matrices, $\bm{C}$ and $\bm{F}^{-1}$, of shape $m\times m$ for which the following inequality holds:
@@
From this, it follows that
@@
$\det\bm{C}$ $\geq\det\bm{F}^{-1}.$ (33)
Here the trick was to push the inequality toward the multivariate Gaussian integral. After integrating, what is left is only the inverse of a normalization constant. We note that if one wants to interpret $\bm{C}$ as a covariance matrix and $\bm{F}$ as a Fisher matrix, then it would make more sense to integrate over $\bm{x}^{T}\bm{C}^{-1}\bm{x}$ and $\bm{x}^{T}\bm{F}\bm{x}$. However, this is irrelevant for the proof, as it would only result in the change of the inequality direction in each row, ending with $\det\bm{C}^{-1}\leq\det\bm{F}$. We usually assume the matrices are nonsingular and thus invertible, which is needed here.
We note that the identical procedure can be used to prove that the one-term Fisher is a lower bound of the two-term Fisher. For $\bm{F}=\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}}$ one can write $\bm{x}^{T}(\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}})\bm{x}\geq\bm{x}^{T}\bm{F}_{\bm{\mu}}\bm{x}$, which by following the same steps translates into $\det(\bm{F}_{\bm{\mu}}+\bm{F}_{\bm{\Sigma}})\geq\det\bm{F}_{\bm{\mu}}$.
@@ Appendix B Derivation of a Gaussian Fisher information matrix
To derive Eq. 4 (i.e., the Fisher information of a multivariate Gaussian distribution), we started by writing the Fisher information matrix equation (Eq. 1) in the equivalent form:
@@
The first one follows from $\bm{\Sigma}\bm{\Sigma}^{-1}=\mathbb{1}$ and calculating its derivative, the second one is Jacobi’s relation for the derivative of the determinant. The last two are expectation relations for Gaussian distributions for an arbitrary matrix $\bm{A}$ (where $\dim\bm{A}=\dim\bm{\Sigma}$). The first of these follows from the antisymmetry of the integral, and the second is left as an exercise for the conscientious reader.
@@
and computing its second derivative, we get
@@
$\phantom{=}-\left(\partial_{m}\bm{\mu}^{T}\bm{\Sigma}^{-1}\partial_{n}\bm{\Sigma}+\partial_{n}\bm{\mu}^{T}\bm{\Sigma}^{-1}\partial_{m}\bm{\Sigma}\right)\bm{\Sigma}^{-1}(\bm{d}-\bm{\mu}).$ (48)
Taking the negative expectation value of the final equation can now be done by using relations 39 and 40. One can easily see that only the first three terms contribute, with others canceling each other out or being equal to zero. With this, we finally have